Grade 12th passMechanics

When resultant bisects the angle between two vectors. Then those vectors have equal magnitudes.
Why?

Ishan karki
8 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago

① THE RESULTANT OF TWO VECTORS OF EQUAL MAGNITUDE

THAT ARE INCLINED AT AN ANGLE θ TO EACH OTHER,bisects θ.

② Let the two vectors be OA,OB ,∠AOB=θ

Let |OA|=|OB|=m

∴Let //gram of forces be AOBR, OR=resultant=r

In △ of forces isosceles △AOR, AR=OB ,

let∠AOR=α,∠OAR=β

α+α+β=180°

β=180-2α↙

But∠OAR=β=180-θ→

↘β=180-θ=180–2α

θ=2α